# LASeR: Learning to Adaptively Select Reward Models with Multi-Armed Bandits

### v2

@@ Appendix C Generalization Capabilities of LASeR
To demonstrate the generalization of LASeR to other Bandit Algorithms, we include empirical comparisons with NeuralUCB, a more expressive contextual bandit and Exp3, a widely used non-contextual bandit algorithm. NeuralUCB is parameterized by a 3-layer MLP with hidden dimensions 256 and 128. The output layer of NeuralUCB produces a single scalar, which is the predicted expected reward for each RM based on the context. Table 9 shows that LinUCB achieves comparable performance to Exp3 and performs slightly below NeuralUCB on 2 out of 3 datasets. However, LinUCB is significantly more efficient and stable because it uses closed-form updates and maintains simple per-arm information, avoiding the overhead of training neural networks as in NeuralUCB. These results support our central claim: regardless of the specific bandit algorithm used, LASeR consistently outperforms baseline RM selection and ensemble baselines.
@@
Generalization to Different Base Models.
In the main paper, we present the results for the Llama-3-8B model. Here we present the results for Mistral-7B model across reasoning, instruction-following and long-context understanding tasks.
@@
In instruction-following tasks, Fig. 10 shows that LASeR consistently outperforms all baselines, including Best RM (58.72%), Sequential RM Selection (63.72%), and Random RM Selection (70.61%). It achieves the highest win rate against RM Score Ensemble (73.27%) and surpasses Online RM Ensemble (65.41%), demonstrating its ability to generalize to different LLMs.
@@
In long-context understanding tasks, Table 2 highlights LASeR’s superior performance on long-context understanding tasks with Mistral-7B. LASeR achieves the highest scores for Single-Doc QA (29.14) and Multi-Doc QA (27.80), significantly outperforming the base model. In Few-Shot Learning, LASeR achieves 68.47, surpassing both the Best RM (68.34) and Online RM Ensemble (66.81). While its Rouge-L score for summarization (30.08) is slightly below Best RM (30.42), LASeR remains comparable. These results demonstrate the consistent performance of LASeR across the model family.
Generalization to the Base and RMs of Bigger Sizes.
To further analyze the generalization of LASeR to the base and RMs of bigger sizes, we add an experiment with Qwen2.5-32B [Qwen Team, 2024a] policy model and the following 4 similarly sized RMs from RewardBench: RISE-Judge-Qwen2.5-32B, Bagel-DPO-34B, Archangel-DPO-30B, Archangel-KTO-30B. We compare LASeR with the best RM and the RM Agreement Ensemble baselines, which are the second-best baselines for most of the benchmarks on Wildchat. Table 12 demonstrates that LASeR has higher win rates than baselines.
@@
Generalization to the Number of RMs.
@@
To study the generalization capability of LASeR across the number of RMs, we expand the candidate set of RMs with up to 4 more RMs from the RewardBench leaderboard, including Tulu-2-DPO-7B [Ivison et al., 2023], Zephyr-7B-Gemma [Tunstall and Schmid, 2024], Qwen1.5-MoE-A2.7B-Chat [Qwen Team, 2024b], Archangel-7B [Ethayarajh et al., 2024]. Fig. 11 shows that the accuracy remains consistent across all datasets as the number of RMs varies. StrategyQA remains near 85.9%, GSM8K around 74.8%, and MMLU close to 68.1%, with minimal fluctuations, indicating robust performance regardless of the number of RMs while RM Agreement Ensemble performance declines more steeply.
Generalization to Training Loss Functions.
@@
In Sec. 3.1, we state that the choice of loss function used to train the LLM depends on the underlying task or domain. Nevertheless, we always use the training loss as the MAB reward to update the MAB’s parameters. Here we study the performance of LASeR and baselines with 4 different loss functions, NLL, DPO, NLL + DPO [Pang et al., 2024], and KTO [Ethayarajh et al., 2024], in the reasoning domain. Results in Table 13 show that training LLMs with multiple rewards using LASeR outperforms sequential RM selection by $2.4\%$, $1.7\%$, and $1.3\%$ when using NLL, DPO, NLL+DPO loss functions, respectively; while both methods yield comparable performance with KTO. Additionally, we found that the most effective training loss functions are NLL + DPO for StrategyQA, NLL for GSM8K, and KTO for MMLU. However, irrespective of the choice of the underlying loss function, LASeR is more effective at balancing and adaptively selecting from multiple RMs. Lastly, we compare LASeR with a variant in which we use $\mathrm{Acc}(y^{w})-\mathrm{Acc}(y^{l})$ as the MAB reward, which uses the ground-truth information about the final answer. We find that using the negative training loss of the LLM is more effective than using accuracy as the MAB reward.
@@
Generalization to OOD Tasks.
We first assess the generalization ability of our method by training models on specific datasets and evaluating their performance on out-of-distribution reasoning tasks. Specifically, we train the model on the StrategyQA and MMLU datasets and evaluate its generalization on the CommonsenseQA [CSQA; Talmor et al., 2019] dataset. Similarly, we train on GSM8K and test on MATH [Hendrycks et al., 2021c] to assess the model’s ability to generalize across different reasoning datasets. These tasks are designed to capture both general reasoning ability and OOD generalization across domains. We report the results in Table 14, where we find that across both Llama-3-8B and Mistral-7B models, models trained with LASeR yield the best average accuracy, beating training with the best RM by roughly $2\%$ (absolute) on CSQA with Llama-3-8B. On Mistral-7B, training with LASeR outperforms both training with single best RM and sequential RM selection by slightly over $1\%$.
@@
Robustness to Noisy Rewards.
To examine the robustness of our method in the presence of noisy or irrelevant rewards, we conduct the following analysis using Llama-3-8B on GSM8K. We add varying amounts of Gaussian noise $\sigma$ to the rewards generated by RMs sampled from the distribution $\mathcal{N}(0,\sigma I)$, where $I$ is the identity matrix, to simulate noisy rewards when using RMs in out-of-distribution settings. In addition to LASeR using the LinUCB algorithm (c.f. Sec. 3.2), we also use Exp3 [Auer et al., 2002] designed for adversarial bandit settings. In Fig. 12, we find that even as the degree of noise in RM scores increases (from $\sigma\!=\!$ 0.1 to 0.4), LASeR’s selection strategy continues to perform robustly, mitigating the effects of noise compared to the sequential baseline. Specifically, LASeR has an average performance drop of only $0.55\%$ accuracy at a noise level of $\sigma\!=\!0.3$, whereas the sequential baseline suffers a $1.6\%$ accuracy drop (3 times as much) under the same conditions. Furthermore, LASeR using Exp3, the most noise-robust method, maintains consistent performance, with only a $0.26\%$ accuracy drop.
@@
LASeR Training with Domain-specific Evaluation Metrics.
While recent works focus on building RMs that reflect preferences across domains, an extensive body of prior work develops a suite of evaluation metrics catered to specific domains such as reasoning [Golovneva et al., 2022, Prasad et al., 2023]. To show that LASeR can be used to select any kind of evaluation metric from a collection of metrics during training, in Table 15, we present results with training LLMs on model-based metrics from ROSCOE [Golovneva et al., 2022] by replacing RMs with informativeness, faithfulness, reasoning alignment, hallucination, common sense error, semantic, coherence and perplexity in Sec. 3. Llama-3-8B models trained using LASeR yield 1.62% accuracy improvement over baselines on average across three datasets. These results are also generalized to Mistral-7B, except for GSM8K, where we achieve comparable performance to the Base + Informativeness baselines. Note that the perplexity of most responses is nearly identical, making it difficult to distinguish between them, explaining why perplexity shows little to no improvement compared to the base model.
@@
Robustness to Underperforming Evaluation Metrics.
Similar to our analysis on noise in rewards in Sec. 5, we investigate how adding ROSCOE metrics with poor correlation to human-annotated labels in meta-evaluation by Golovneva et al. [2022] impacts the performance of Llama-3-8B on GSM8K. Once again, even with ROSCOE metrics, demonstrates LASeR can maintain consistent performance by adaptively prioritizing the most relevant reward signals, outperforming the sequential baseline, which fails to filter out irrelevant information effectively. Fig. 13 shows that as the number of irrelevant metrics increases, LASeR’s selection strategy continues to perform robustly. Specifically, LASeR has an average performance drop of only $0.13\%$, whereas the sequential baseline suffers a $2.15\%$ accuracy drop under the same conditions. Lastly, LASeR using Exp3 maintains a consistent performance level with a $0.4\%$ accuracy drop.
@@
Appendix D Additional Discussion
@@
LASeR with different “kinds” of RMs. In Sec. 4.2, we show that LASeR can choose from a set of candidate RMs, and our analysis in Fig. 12 highlights the fact that LASeR is robust to noisy RMs. In Appendix C, we show that LASeR can also be used with metric-based rewards [Golovneva et al., 2022]. These experiments reflect a conceptual split between the generator (the LLM) and the scorer (the RM or metric). Thus, LASeR is applicable to other settings that follow this paradigm, e.g., using an LLM-as-a-judge [Zheng et al., 2023], where LASeR could be used to choose between different judge models, prompts, or different combinations of RMs and metrics. However, consistent with the “self-preference” bias of LLMs [Panickssery et al., 2024], we caution that using an RM that is based on the same model as the LLM used for generating responses could lead to the MAB spuriously favoring certain RMs. We leave further study on extending LASeR to future work.
@@
Quality of RMs used with LASeR. Methods that rely on RMs for scoring generally assume that these RMs have a strong correlation with human judgments. LASeR tempers this assumption in a number of ways: First, by ensembling multiple RMs, LASeR weakens the effect of noisy RMs; this can be seen in Fig. 12, where LASeR mitigates the negative impact of a noisy RM even as the level of noise is increased. Moreover, the fact that LASeR can select RMs at an instance level means that there need not be a single RM that always correlates well across all instances. However, LASeR does require at least one RM to be positively correlated with human judgments on each instance. If this assumption is not met (i.e., all RMs are poorly correlated across all instances), then optimizing for the RMs will yield poor results. Note that this holds true for any method optimizing for RMs. Because LASeR selects from multiple RMs, its contributions are complementary to developments in RMs, which can easily be integrated into LASeR, as well as improvements to preference optimization loss functions (see Appendix C). Such RM improvements are likely to be necessary as LLMs are deployed in domains that are out of scope for existing systems and domains with heterogeneous requirements (e.g., our generation domains in Sec. 4.2). In these cases, there will be no single “perfect” existing RM, and successful solutions will likely involve mixing multiple RMs. A core benefit of LASeR is its ability to automatically filter RMs; in Fig. 6 we see that utilization differs across domains. This allows users to avoid expensive experimentation with subsets of RMs: they can simply offload this task to LASeR, which will automatically select the more useful RM(s).
Appendix E Prompts

### v3

@@ Appendix C Generalization Capabilities of LASeR
To demonstrate the generalization of LASeR to other Bandit Algorithms, we include empirical comparisons with NeuralUCB, a more expressive contextual bandit and Exp3, a widely used non-contextual bandit algorithm. NeuralUCB is parameterized by a 3-layer MLP with hidden dimensions 256 and 128. The output layer of NeuralUCB produces a single scalar, which is the predicted expected reward for each RM based on the context. Table˜11 shows that LinUCB achieves comparable performance to Exp3 and performs slightly below NeuralUCB on 2 out of 3 datasets. However, LinUCB is significantly more efficient and stable because it uses closed-form updates and maintains simple per-arm information, avoiding the overhead of training neural networks as in NeuralUCB. These results support our central claim: regardless of the specific bandit algorithm used, LASeR consistently outperforms baseline RM selection and ensemble baselines.
@@
Generalization to Different Base Models.  In the main paper, we present the results for the Llama-3-8B model. Here we present the results for Mistral-7B model across reasoning, instruction-following and long-context understanding tasks.
@@
In instruction-following tasks, Fig.˜10 shows that LASeR consistently outperforms all baselines, including Best RM (58.72%), Sequential RM Selection (63.72%), and Random RM Selection (70.61%). It achieves the highest win rate against RM Score Ensemble (73.27%) and surpasses Online RM Ensemble (65.41%), demonstrating its ability to generalize to different LLMs.
@@
In long-context understanding tasks, Table˜2 highlights LASeR’s superior performance on long-context understanding tasks with Mistral-7B. LASeR achieves the highest scores for Single-Doc QA (29.14) and Multi-Doc QA (27.80), significantly outperforming the base model. In Few-Shot Learning, LASeR achieves 68.47, surpassing both the Best RM (68.34) and Online RM Ensemble (66.81). While its Rouge-L score for summarization (30.08) is slightly below Best RM (30.42), LASeR remains comparable. These results demonstrate the consistent performance of LASeR across the model family.
Generalization to the Base and RMs of Bigger Sizes.  To further analyze the generalization of LASeR to the base and RMs of bigger sizes, we add an experiment with Qwen2.5-32B [Qwen Team, 2024a] policy model and the following 4 similarly sized RMs from RewardBench: RISE-Judge-Qwen2.5-32B, Bagel-DPO-34B, Archangel-DPO-30B, Archangel-KTO-30B. We compare LASeR with the best RM and the RM Agreement Ensemble baselines, which are the second-best baselines for most of the benchmarks on Wildchat. Table˜14 demonstrates that LASeR has higher win rates than baselines.
@@
Generalization to Correctness Rewards for Reasoning.  As verifiable rewards emerge as a powerful RM for reasoning tasks, in this experiment, we analyze how does it compared with current multiple RMs framework of LASeR, and whether it is possible to incorporate them into LASeR.
First, we implement a Correctness Reward baseline for the reasoning datasets. Specifically, we assign a reward of 1 if the generated answer matches the correct final answer and 0 otherwise. We maintain the same iterative training setup as in our main experiments and compare with LASeR training with multiple RMs. Table˜15 shows that while the correctness-based reward performs competitively, especially on GSM8K (74.56% vs. 74.75% accuracy), LASeR outperforms it across all tasks, including commonsense reasoning tasks (StrategyQA and MMLU). This is likely because correctness-based rewards are well-suited for tasks with clearly verifiable outputs, such as math or code, but they struggle to capture the multi-step reasoning and contextual understanding needed in factual or open-domain QA, where answers are often ambiguous or depend on background knowledge. In these cases, other forms of supervision such as process-based rewards [Ma et al., 2025] or human preference signals modeled by RMs are typically more effective. LASeR is designed to leverage this by selecting among multiple RMs, allowing it to perform well across a variety of task types.
Table 15: Performance comparison of LASeR and the Correctness Reward baseline.
Method StrategyQA GSM8K MMLU Avg.
Correctness Reward 83.42 74.56 66.35 74.78
LASeR 85.96 74.75 68.24 76.32
As noted in Sec.˜3, LASeR can treat any RMs as part of its selection framework. We show that LASeR can integrate correctness-based rewards, combining their strengths with those of preference-based models. To validate that, we run an additional experiment on GSM8K, where we add the correctness-based RM to the existing pool of learned RMs (Olmo, Eurus, Qwen, Zephyr). To further test LASeR’s ability to adaptively select appropriate RMs, we create a mixed dataset combining GSM8K with WildChat (subsampled to 1.2K examples per category to ensure balance). Since correctness-based signals are not applicable to WildChat’s open-ended prompts, we assign a random binary reward (0 or 1) for those examples to simulate a misleading RM. We observe that LASeR achieves an accuracy of 75.06% accuracy on the GSM8K test set, improving over LASeR without correctness rewards by 0.31%. We also report the utilization rate of each RM in the pool (Olmo, Eurus, Qwen, Zephyr, correctness/random) for two types of queries: GSM8K and WildChat in Table˜16. On GSM8K, LASeR heavily leverages the correctness-based RM (40.26%) alongside strong learned RMs like Qwen (43.58%), indicating that it effectively identifies and uses the most reliable supervision. In contrast, for WildChat, LASeR downweights the misleading correctness-based RM (2.84%) and shifts toward more suitable RMs like Zephyr and Qwen. These results demonstrate LASeR’s ability to adapt RM usage based on task domain, even in the presence of noisy or misleading supervision.
Table 16: Utilization rate (%) of different RMs including a correctness RM and a random RM on GSM8K and WildChat.
Utilization Rate (%) Olmo Eurus Qwen Zephyr Correctness/Random
GSM8K 4.02 2.32 43.58 19.82 40.26
WildChat 9.09 16.41 27.75 45.90 2.84
@@
Generalization to the Number of RMs.  To study the generalization capability of LASeR across the number of RMs, we expand the candidate set of RMs with up to 4 more RMs from the RewardBench leaderboard, including Tulu-2-DPO-7B [Ivison et al., 2023], Zephyr-7B-Gemma [Tunstall and Schmid, 2024], Qwen1.5-MoE-A2.7B-Chat [Qwen Team, 2024b], Archangel-7B [Ethayarajh et al., 2024]. Fig.˜11 shows that the accuracy remains consistent across all datasets as the number of RMs varies. StrategyQA remains near 85.9%, GSM8K around 74.8%, and MMLU close to 68.1%, with minimal fluctuations, indicating robust performance regardless of the number of RMs while RM Agreement Ensemble performance declines more steeply.
@@
Generalization to Training Loss Functions.  In Sec.˜3.1, we state that the choice of loss function used to train the LLM depends on the underlying task or domain. Nevertheless, we always use the training loss as the MAB reward to update the MAB’s parameters. Here we study the performance of LASeR and baselines with 4 different loss functions, NLL, DPO, NLL + DPO [Pang et al., 2024], and KTO [Ethayarajh et al., 2024], in the reasoning domain. Results in Table˜17 show that training LLMs with multiple rewards using LASeR outperforms sequential RM selection by $2.4\%$, $1.7\%$, and $1.3\%$ when using NLL, DPO, NLL+DPO loss functions, respectively; while both methods yield comparable performance with KTO. Additionally, we found that the most effective training loss functions are NLL + DPO for StrategyQA, NLL for GSM8K, and KTO for MMLU. However, irrespective of the choice of the underlying loss function, LASeR is more effective at balancing and adaptively selecting from multiple RMs. Lastly, we compare LASeR with a variant in which we use $\mathrm{Acc}(y^{w})-\mathrm{Acc}(y^{l})$ as the MAB reward, which uses the ground-truth information about the final answer. We find that using the negative training loss of the LLM is more effective than using accuracy as the MAB reward.
@@
Generalization to OOD Tasks.  We first assess the generalization ability of our method by training models on specific datasets and evaluating their performance on out-of-distribution reasoning tasks. Specifically, we train the model on the StrategyQA and MMLU datasets and evaluate its generalization on the CommonsenseQA [CSQA; Talmor et al., 2019] dataset. Similarly, we train on GSM8K and test on MATH [Hendrycks et al., 2021c] to assess the model’s ability to generalize across different reasoning datasets. These tasks are designed to capture both general reasoning ability and OOD generalization across domains. We report the results in Table 18, where we find that across both Llama-3-8B and Mistral-7B models, models trained with LASeR yield the best average accuracy, beating training with the best RM by roughly $2\%$ (absolute) on CSQA with Llama-3-8B. On Mistral-7B, training with LASeR outperforms both training with single best RM and sequential RM selection by slightly over $1\%$.
@@
Robustness to Noisy Rewards.  To examine the robustness of our method in the presence of noisy or irrelevant rewards, we conduct the following analysis using Llama-3-8B on GSM8K. We add varying amounts of Gaussian noise $\sigma$ to the rewards generated by RMs sampled from the distribution $\mathcal{N}(0,\sigma I)$, where $I$ is the identity matrix, to simulate noisy rewards when using RMs in out-of-distribution settings. In addition to LASeR using the LinUCB algorithm (c.f. Sec.˜3.2), we also use Exp3 [Auer et al., 2002] designed for adversarial bandit settings. In Fig.˜12, we find that even as the degree of noise in RM scores increases (from $\sigma\!=\!$ 0.1 to 0.4), LASeR’s selection strategy continues to perform robustly, mitigating the effects of noise compared to the sequential baseline. Specifically, LASeR has an average performance drop of only $0.55\%$ accuracy at a noise level of $\sigma\!=\!0.3$, whereas the sequential baseline suffers a $1.6\%$ accuracy drop (3 times as much) under the same conditions. Furthermore, LASeR using Exp3, the most noise-robust method, maintains consistent performance, with only a $0.26\%$ accuracy drop.
@@
LASeR Training with Domain-specific Evaluation Metrics.  While recent works focus on building RMs that reflect preferences across domains, an extensive body of prior work develops a suite of evaluation metrics catered to specific domains such as reasoning [Golovneva et al., 2022, Prasad et al., 2023]. To show that LASeR can be used to select any kind of evaluation metric from a collection of metrics during training, in Table 19, we present results with training LLMs on model-based metrics from ROSCOE [Golovneva et al., 2022] by replacing RMs with informativeness, faithfulness, reasoning alignment, hallucination, common sense error, semantic, coherence and perplexity in Sec.˜3. Llama-3-8B models trained using LASeR yield 1.62% accuracy improvement over baselines on average across three datasets. These results are also generalized to Mistral-7B, except for GSM8K, where we achieve comparable performance to the Base + Informativeness baselines. Note that the perplexity of most responses is nearly identical, making it difficult to distinguish between them, explaining why perplexity shows little to no improvement compared to the base model.
@@
Robustness to Underperforming Evaluation Metrics.  Similar to our analysis on noise in rewards in Sec.˜5, we investigate how adding ROSCOE metrics with poor correlation to human-annotated labels in meta-evaluation by Golovneva et al. [2022] impacts the performance of Llama-3-8B on GSM8K. Once again, even with ROSCOE metrics, demonstrates LASeR can maintain consistent performance by adaptively prioritizing the most relevant reward signals, outperforming the sequential baseline, which fails to filter out irrelevant information effectively. Fig.˜13 shows that as the number of irrelevant metrics increases, LASeR’s selection strategy continues to perform robustly. Specifically, LASeR has an average performance drop of only $0.13\%$, whereas the sequential baseline suffers a $2.15\%$ accuracy drop under the same conditions. Lastly, LASeR using Exp3 maintains a consistent performance level with a $0.4\%$ accuracy drop.
@@
Appendix D Ablation Study
Context Embedding.  As noted in Sec.˜3, to validate the effectiveness of the last-token embedding method, we conduct an ablation study comparing three context embedding approaches: first-token (using the embedding of the first token in the prompt), average-token (using the average of all token embeddings in the prompt), and last-token (using the embedding of the last token in the prompt). We evaluate these methods on Llama-3-8B and reasoning tasks. Table˜20 shows that the last-token embeddings consistently outperform the alternatives across all three datasets, demonstrating the effectiveness of this method in LASeR.
Table 20: Comparing different token embedding methods in LASeR.
Method StrategyQA GSM8K MMLU Avg.
LASeR (first) 83.59 73.03 67.03 74.55
LASeR (avg.) 83.22 73.47 66.52 74.40
LASeR (last) 85.96 74.75 68.24 76.32
Impact of Bandit Parameter $\alpha$.  We conduct an ablation study to assess the impact of the exploration parameter $\alpha$ in the LinUCB algorithm on Llama-3-8B and reasoning tasks. Table˜21 shows that for reasoning datasets, disabling exploration leads to suboptimal performance, particularly in early training stages where the bandit may overfit to a suboptimal RM. In contrast, enabling exploration ($\alpha>0$) improves performance by allowing the model to discover and leverage more informative RMs over time. We find that setting $\alpha$ in the range of 0.4 to 0.6 consistently yields the best results across tasks, offering a good balance between exploration and exploitation. This is consistent with our observations on the performance on dev set for reasoning tasks as noted in Sec.˜A.1.
Table 21: Effect of varying $\alpha$ on accuracy for reasoning tasks.
Dataset $\alpha$=0.0 $\alpha$=0.2 $\alpha$=0.4 $\alpha$=0.6 $\alpha$=0.8 $\alpha$=1.0
StrategyQA 83.11 85.29 85.96 85.32 84.53 83.77
GSM8K 71.68 73.22 74.10 74.75 74.63 73.94
Appendix E Additional Discussion
@@
LASeR with different “kinds” of RMs.  In Sec.˜4.2, we show that LASeR can choose from a set of candidate RMs, and our analysis in Fig.˜12 highlights the fact that LASeR is robust to noisy RMs. In Appendix˜C, we show that LASeR can also be used with metric-based rewards [Golovneva et al., 2022]. These experiments reflect a conceptual split between the generator (the LLM) and the scorer (the RM or metric). Thus, LASeR is applicable to other settings that follow this paradigm, e.g., using an LLM-as-a-judge [Zheng et al., 2023], where LASeR could be used to choose between different judge models, prompts, or different combinations of RMs and metrics. However, consistent with the “self-preference” bias of LLMs [Panickssery et al., 2024], we caution that using an RM that is based on the same model as the LLM used for generating responses could lead to the MAB spuriously favoring certain RMs. We leave further study on extending LASeR to future work.
@@
Quality of RMs used with LASeR.  Methods that rely on RMs for scoring generally assume that these RMs have a strong correlation with human judgments. LASeR tempers this assumption in a number of ways: First, by ensembling multiple RMs, LASeR weakens the effect of noisy RMs; this can be seen in Fig.˜12, where LASeR mitigates the negative impact of a noisy RM even as the level of noise is increased. Moreover, the fact that LASeR can select RMs at an instance level means that there need not be a single RM that always correlates well across all instances. However, LASeR does require at least one RM to be positively correlated with human judgments on each instance. If this assumption is not met (i.e., all RMs are poorly correlated across all instances), then optimizing for the RMs will yield poor results. Note that this holds true for any method optimizing for RMs. Because LASeR selects from multiple RMs, its contributions are complementary to developments in RMs, which can easily be integrated into LASeR, as well as improvements to preference optimization loss functions (see Appendix˜C). Such RM improvements are likely to be necessary as LLMs are deployed in domains that are out of scope for existing systems and domains with heterogeneous requirements (e.g., our generation domains in Sec.˜4.2). In these cases, there will be no single “perfect” existing RM, and successful solutions will likely involve mixing multiple RMs. A core benefit of LASeR is its ability to automatically filter RMs; in Fig.˜6 we see that utilization differs across domains. This allows users to avoid expensive experimentation with subsets of RMs: they can simply offload this task to LASeR, which will automatically select the more useful RM(s).
Appendix F Prompts